OpenStudy (anonymous):

decompose into partial fractions (x^2+x-5)/((x-1)(x^2-2))

OpenStudy (anonymous):

Lagrangeson!

OpenStudy (anonymous):

Ax +b C ------ + ------- x^2-2 (x-1)

OpenStudy (anonymous):

why ax+b?

OpenStudy (anonymous):

because (x^2-2) is not factorable into linear

OpenStudy (anonymous):

It's \(ax+b\) because the denominator of that fraction has degree 2. You can also bypass this by decomposing in the following manner.\[\frac{x^2+x-5}{(x-1)(x^2-2)} = \frac{A}{x+\sqrt 2}+\frac{B}{x-\sqrt{2}}+\frac{C}{x-1}\]

OpenStudy (anonymous):

oh, thx